sir, how to solve --  2^log18tobase6 * 3^log3tobase6


 


 


 


 


 


 


 


 


 


 

Grade:11

2 Answers

Sagar Honnungar
13 years ago

Since log 6 18= log 6 6*3=log 6 6 + log 6 3=1+log 6 3 ,

 the question reduces to 21+log63*3log63= 2*6log63 = 2*3 = 6

rhea gupta
13 years ago

QUESTION:

2^log18tobase6 * 3^log3tobase6

ANSWER:

As  log 6 18= log 6*3=log 6 6 + log 3=1+log 

Therefore, 21+log63*3log63= 2*6log63 = 2*3 = 6

( BASICALLY IT REQUIRES PROPERTIES OF LOGARITMS)